The reactions of hydrocarbons are often studied in the petroleum industry. One such reaction is 2C3H8(g) → C6H6(ℓ) + 5H2(g) with ΔH° = 698 kJ. If 30 L of propane (C3H8) at 25°C and 0.90 atm is to react, how much heat must be supplied (in KJ)? R = 0.08214 L atm mol-1 K-1
Answer #1

Volume (V) = 30 L

Pressure (P) = 0.90 atm

Temperature (T) =25° C = 298 K

Let n is moles

Using ideal gas equation;

PV = nRT

n = PV/RT = 0.90 * 30 / 0.08214 * 298 = 0.8259 moles of propane

From reaction;

2 mole C3H8 requires = 698 KJ

So, 0.8259 mole will require = 698*0.8259/2 = 288.24 KJ heat

So, 288.24 KJ heat must be supplied.